The larger of two numbers is 6 less than twice the smaller number. If their total is 27, find both numbers
Answer by lenny460(1073)

The larger of two numbers is 6 less than twice the smaller number. If their total is 27, find both numbers ?

Let the larger number be x.
Let the smaller number be y
Therefore:
x + y = 27
y = 27 - x
x = 2y - 6
x = 2(27 - x) - 6
x = 54 - 2x - 6
54 - 6 = 48
x = 48 - 2x
Add 2x to each side of the equation.
x + 2x = 48 - 2x + 2x
3x = 48
x = 16
y = 27 - 16
y = 11

Answer:
x = 16 and
y = 11



Lennox Obuong
Algebra Tutor
Nairobi, Kenya